The sum of three consecutive integers is 177. What is the largest of the three numbers?


Answer by Maths68(1474)

Let x+1, x+2 and x+3 be the three consecutive numbers
Given
The sum of three consecutive integers is 177
x+1+x+2+x+3=177
3x+6=177
3x=177-6
3x=171
3x/3=171/3
x=57


Numbers are
x+1=57+1=58
x+2=57+2=59
x+3=57+3=60
The largest of the three numbers is 60.